OpenStudy (anonymous):

y2-y1/x2-x1=??? when x1=24,, x2=-27,,y1=18,,y2=12 and it needs to be simplfied answer

OpenStudy (anonymous):

\[\frac{y_2-y_2}{x_2-x_1}\]

OpenStudy (anonymous):

in this case you get \[\frac{-27-24}{12-18}\]

OpenStudy (anonymous):

better known as \[\frac{-51}{-6}=\frac{17}{2}=8.5\]

OpenStudy (lalaly):

noo its y2-y1/x2-x1 so its 2/17

OpenStudy (anonymous):

-6/51

OpenStudy (anonymous):

what did i do it upside down?

OpenStudy (lalaly):

lol yeh

OpenStudy (anonymous):

oh damn sorry

OpenStudy (anonymous):

thanks everyone this is alot of help

OpenStudy (anonymous):

thanks everyone this is alot of help

OpenStudy (anonymous):

then flip it \[\frac{12-18}{-27-24}=\frac{-6}{-51}=\frac{2}{17}\]